OpenStudy (anonymous):

Differentiate f(x)=Coshx/Sinh(x2) can anybody give me the solution to this? Thank you...

OpenStudy (anonymous):

thats is x square.....

terenzreignz (terenzreignz):

Well, basics... quotient rule, first.

terenzreignz (terenzreignz):

\[\large \frac{d}{dx}\frac{f(x)}{g(x)}=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\]

terenzreignz (terenzreignz):

In this case f(x)=cosh x g(x)= sinh x²

OpenStudy (anonymous):

yes I did that... but the answer am getting is not correct... I think it requires trigonometric substitutions...

terenzreignz (terenzreignz):

Could you post your solution?

OpenStudy (anonymous):

Yes one minute...

OpenStudy (anonymous):

\[Sinhx ^{2} .Sinhx+ Coshx. 2x Coshx ^{2} / (Sinhx ^{2})^{2}\]

terenzreignz (terenzreignz):

I for one, don't see what's wrong with it...

OpenStudy (anonymous):

how to further simplify that?

terenzreignz (terenzreignz):

None that I know of... Hyperbolic functions confuse me somewhat... unless you're suggesting we expand the sinh and cosh into the forms involving powers of e.

OpenStudy (anonymous):

ohh then I guess this is the final answer.....

OpenStudy (sirm3d):

shouldn't it be ( ) ( ) MINUS ( ) ( )

terenzreignz (terenzreignz):

!!!!!!!! Completely overlooked that. Bloody details -.- It should indeed be sinh x² sinh x + 2x cosh x cosh x² in the numerator

OpenStudy (anonymous):

no no it is PLUS...

terenzreignz (terenzreignz):

My bad.

terenzreignz (terenzreignz):

It's not plus... unlike the trigonometric cos the derivative of cosh x is sinh x

terenzreignz (terenzreignz):

^as opposed to the derivative of cos x being -sin x

OpenStudy (anonymous):

well, it is my fist day here, did not expect people would respond, am amazed and very happy I found this site... thank you :) :) @terenzreignz @sirm3d ..... I guess it cannot be simplified any further in trig terms... thank you :)

terenzreignz (terenzreignz):

No problem :)

OpenStudy (sirm3d):

you're welcome.